# From LLMs to MLLMs to Agents: A Survey of Emerging Paradigms in Jailbreak Attacks and Defenses within LLM Ecosystem

### v2

@@ Abstract
Large language models (LLMs) are rapidly evolving from single-modal systems to multimodal LLMs and intelligent agents, significantly expanding their capabilities while introducing increasingly severe security risks. This paper presents a systematic survey of the growing complexity of jailbreak attacks and corresponding defense mechanisms within the expanding LLM ecosystem. We first trace the developmental trajectory from LLMs to MLLMs and Agents, highlighting the core security challenges emerging at each stage. Next, we categorize mainstream jailbreak techniques from both the attack impact and visibility perspectives, and provide a comprehensive analysis of representative attack methods, related datasets, and evaluation metrics. On the defense side, we organize existing strategies based on response timing and technical approach, offering a structured understanding of their applicability and implementation. Furthermore, we identify key limitations in existing surveys, such as insufficient attention to agent-specific security issues, the absence of a clear taxonomy for hybrid jailbreak methods, a lack of detailed analysis of experimental setups, and outdated coverage of recent advancements. To address these limitations, we provide an updated synthesis of recent work and outline future research directions in areas such as dataset construction, evaluation framework optimization, and strategy generalization. Our study seeks to enhance the understanding of jailbreak mechanisms and facilitate the advancement of more resilient and adaptive defense strategies in the context of ever more capable LLMs.
@@
LLMs, MLLMs, agents, jailbreak attack, defense strategy
@@ I-A Development of LLMs
The evolution of neural network architectures undergoes multiple paradigm shifts. Early sequence modeling primarily relies on Recurrent Neural Networks (RNNs [1]), whose performance is limited by the vanishing gradient problem when modeling long-term dependencies. Long Short-Term Memory networks (LSTMs [2]) alleviate this issue to some extent by introducing various gating mechanisms. However, due to their sequential nature, LSTMs remain computationally inefficient when processing large-scale data. The emergence of the Transformer [3] architecture in 2017 fundamentally changes this landscape. Its self-attention mechanism enables global context modeling and supports parallel computation. In the field of computer vision, Convolutional Neural Networks (CNNs [4]) long dominate image processing tasks due to their local receptive fields. The introduction of techniques such as Residual Networks (ResNet [5]) and Layer Normalization [6] makes it feasible to train ultra-deep networks, laying the foundation for large-scale models. Meanwhile, the application of Transformers to vision tasks demonstrates the advantages of global attention mechanisms.
The rapid growth in model size is driven by the scaling laws. The performance-compute power law [7] proposed by OpenAI in 2020 shows that model performance improves with increasing model parameters, data volume, and compute resources. The parameter count increases dramatically in just a few years: from BERT’s [8] 340 million parameters in 2018, to GPT-3’s [9] 175 billion in 2020, and to PaLM’s [10] 540 billion in 2022. Multimodal extensions also become a key focus. CLIP [11] achieves semantic alignment between images and texts through contrastive learning, while ViT [12] validates the Transformer’s feasibility for vision tasks.
In recent years, large language models (LLMs) enter a new phase marked by emergent intelligence. Once model parameters exceed a certain threshold, capabilities such as chain-of-thought reasoning and in-context learning emerge. Technological pathways also diversify: InstructGPT [13] enhances alignment with human intent through instruction tuning; reinforcement learning from human feedback (RLHF [13]) becomes central to value alignment; and parameter-efficient fine-tuning methods such as LoRA [14] reduce adaptation costs. Multimodal integration also accelerates: GPT-4V [15] supports both visual and textual understanding and generation. Recently, DeepSeek-R1 [16] achieves a breakthrough in the LLM field. Based on an improved Mixture of Experts (MoE) architecture [17, 18], it employs dynamic routing for expert load balancing and integrates a shared attention backbone, achieving both computational efficiency and strong generalization across tasks. It demonstrates competitive performance in long-context modeling and complex reasoning tasks.
@@ I-B Advances in LLM Jailbreaking and Defense
With the rapid development of LLMs, their application scenarios expand from pure text processing to multimodal understanding and autonomous agents, reflecting an evolutionary trend from LLMs to Multimodal LLMs (MLLMs) and further to intelligent agents (Agents) [19, 20]. This evolution greatly expands the capabilities of such models, enabling them to handle more complex tasks, but it also introduces more severe security challenges [21, 22]. Among these, jailbreak attacks aim at bypassing safety mechanisms and inducing models to produce inappropriate or restricted content, and they become increasingly complex and diverse [23, 24].
@@
Nevertheless, current jailbreak and defense surveys in LLM ecosystems face several limitations: (1) Insufficient focus on agents: Despite the rapid advancement of agent technology, current jailbreak studies predominantly target traditional LLMs. There is a lack of systematic analysis on jailbreak attacks, adversarial strategies, and defenses specific to agents, which hinders a comprehensive understanding of agent security and limits the optimization of defense mechanisms. (2) Inadequate taxonomy of hybrid jailbreak methods: Modern jailbreak techniques evolve into hybrid forms that combine multiple strategies. These often share common core modules, requiring researchers to carefully classify and analyze them to reveal their structural and developmental logic. (3) Lack of detailed analysis of experimental settings: While various datasets and evaluation metrics are used in jailbreak research, existing surveys rarely provide comprehensive mapping between methods and their evaluation frameworks, impeding comparability and obscuring the strengths and weaknesses of different approaches. (4) Difficulty in covering the latest advancements: Due to the rapid progress of jailbreak and defense research, existing reviews may not incorporate the most recent techniques, resulting in outdated insights into current trends.
To address these limitations and enhance the understanding of jailbreak and defense across the LLM ecosystem, a more comprehensive and up-to-date survey is essential. In this study, we systematically review the evolution from LLMs to MLLMs and then to Agents, analyzing core technologies, key features, and security challenges at each stage. We also categorize mainstream jailbreak methods from both impact of attack and visibility of attack perspectives, along with a detailed summary of associated datasets and evaluation metrics. Moreover, we classify existing defense strategies by their response timing and technical approaches. Finally, we outline open research problems, including dataset construction, diversified attack strategies, and evaluation framework optimization, aiming to provide valuable guidance for future research. Figure 1 illustrates the overall framework of this work, which facilitates a deeper understanding of recent advancements and promotes the development of more effective jailbreak and defense mechanisms.
Figure 1: This paper presents a comprehensive analytical architecture.
@@
(1) We provide a systematic review of the evolution from LLMs to MLLMs and then to Agents, highlighting the task definitions and key features of jailbreak attacks at each stage (Chapter II). By analyzing capability improvements, expanded applications, and emerging security challenges, we offer a structured background for understanding jailbreak and defense mechanisms.
(2) We categorize jailbreak techniques from two perspectives: the impact of attack and the visibility of attack (Chapter III). From the perspective of attack impact, we classify existing methods based on the stage of impact and the hierarchy of impact (Section III-A). From the perspective of attack visibility, we divide attacks into black-box and white-box categories, and further organize them according to their targets (Section III-B).
@@
(4) We classify existing defense strategies according to response timing and technical approach (Chapter VI). The response timing includes input-level, output-level, and joint defenses (Section VI-A), while technical approaches are grouped into rule/heuristic-based, ML/DL-based, adversarial detection, and hybrid strategies (Section VI-B).
@@
Figure 2: Attack workflow of traditional large language models under the text modality. Attacker strategically design adversarial inputs in the textual modality to elicit harmful responses from the model.
Figure 3: Attack workflow of multimodal large language models. Attacker crafts adversarial inputs by exploiting the vulnerabilities across different modalities in combination, aiming to manipulate the model into generating harmful outputs.
@@ II-A From LLMs to MLLMs to agents
LLMs are primarily trained on massive text corpora and focus on capabilities in text generation and comprehension. Through self-attention mechanisms, they capture linguistic patterns and can perform tasks such as question answering, writing, and translation [55]. Despite their impressive performance in text-based interactions, LLMs may still produce erroneous or inappropriate content due to biases in training data or prompt manipulation. Therefore, techniques like safety filtering and reinforcement learning with human feedback (RLHF) [13] are necessary to ensure security. Compared to MLLMs, LLMs are limited to pure text interaction and lack the ability to process multimodal information such as images or audio.
MLLMs overcome the limitations of single-modality text input by integrating visual, auditory, and other types of data, enabling cross-modal reasoning and generation (e.g., storytelling based on images, summarizing video content) [11]. Their core technologies include cross-modal alignment and joint representation learning, which equip them with richer perceptual and expressive capabilities [56]. However, they still face challenges such as noise in multimodal data alignment, implicit semantic conflicts, and increased computational demands. Approaches like dynamic modality weighting and adversarial training are needed to improve cross-modal consistency. Compared to Agents, MLLMs still rely on human instructions for decision-making and lack autonomous action capabilities.
Agents consist of four key components: core, planning, tools, and memory [41, 57]. The core of an LLM-based agent is the LLM itself, while tools refer to external applications and software interfaces that the LLM can invoke, such as internet search, database retrieval, and external system control. The use of tools enables real-time and flexible generation. The planning component is designed to mitigate hallucinations and inaccuracies in the LLM’s outputs, typically through structured prompts and the integration of additional logical frameworks to guide the core model in making accurate decisions. The memory component is another crucial part of LLM agents, addressing the context length limitations of current LLMs by effectively managing and storing large volumes of information. It serves not only as a data repository but also incorporates necessary details in ongoing interactions, ensuring the LLM has access to all relevant information.
@@ II-B Task Definition
As illustrated in Figure 2, jailbreak attacks on LLMs represent a security threat targeting advanced natural language processing systems. Attackers craft prompts or input sequences to bypass the model’s safety mechanisms, thereby inducing the model to generate content that violates ethical guidelines or contains harmful information [58]. Such attacks can be formalized as an optimization problem: given a target output $y_{attack}$, the attacker seeks the optimal input $x^{*}$ such that
@@
where $M$ denotes the target language model and $P$ represents the model’s probability distribution function. This optimization process aims to maximize the likelihood that the model generates the target adversarial output [59].
@@
$\mathbf{x_{multi}^{*}}=\underset{\mathbf{x_{multi}}}{\text{argmax}}\,P(M(\mathbf{x_{multi}})=y_{attack})$ (2)
where $\mathbf{x_{multi}}=\{x_{text},x_{image}\}$ represents the multimodal input features, and $M$ is the multimodal model’s joint representation function. This type of attack not only increases stealth but also raises the complexity of defense mechanisms [62].
As depicted in Figure 4, jailbreak attacks on intelligent agents (Agents) exhibit distinct characteristics. Their core objective is to alter the agent’s decision-making behavior, causing it to deviate from its predefined objective function [63, 64]. From the perspective of reinforcement learning, such attacks can be executed by manipulating the reward function $R(s)$ or the state transition function $T(s,a)$. Specifically, the attacker seeks an optimal policy $\pi^{*}$ such that
@@
where $\gamma$ is the discount factor. Through this optimization, the agent $A$ is induced to select unintended actions $a_{attack}$, defined as:
@@
where $Q$ denotes the action-value function.
Although the aforementioned three types of jailbreak attacks differ in terms of targets, implementation methods, and attack intentions, they all follow a common pattern of achieving adversarial goals through specific inputs or environmental perturbations. From the perspective of technical complexity, LLM jailbreak attacks primarily focus on the generation of adversarial examples at the textual level [65], MLLM jailbreaks involve joint optimization of multimodal features [66], while attacks on Agents require a deep understanding of the agent’s decision-making mechanisms, task planning, and execution frameworks [67]. The evolution of these attack methods not only reflects the increasing complexity of security threats to AI systems but also underscores the importance of systematically organizing and categorizing these attack strategies.
@@ III Jailbreak Methods
In this section, we focus on various jailbreak attack methods from two distinct perspectives: the impact perspective and the visibility perspective. From the impact perspective, we categorize all methods based on the impact stages and the impact hierarchy. From the visibility perspective, we classify jailbreaks into black-box and white-box attacks, and further organize them according to their targets, such as LLMs, MLLMs, and Agents.
@@ III-A Impact of Attack
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=root-line-color, rounded corners, align=center, text centered, minimum width=10em, edge path= [draw=edge-color, line width=1pt] (!parent.east) – +(5pt,0) —- (.child anchor); , s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1pt, par/.style=rotate=0, child anchor=north, parent anchor=south, anchor=center, minimum height=3em, , where level=1 draw=child-one-line-color, text width=15em, minimum height=3em, font=, , where level=2 draw=child-two-line-color, text width=11em, minimum height=3em, font=, , where level=3 draw=child-three-line-color, minimum width=12em, minimum height=3em, font=, , where level=4 draw=child-four-line-color, minimum height=3em, font=, , where level=5 draw=child-five-line-color, minimum height=3em, font=, , [ Impact of Attack, par, for tree=fill=root-color, [ Stages Based on IoA §$\S$ III-A, for tree=fill=child-one-color, [ Training Stage III-A1, for tree=fill=child-two-color, [ Backdoor Attack, for tree=fill=child-three-color [ [68] [69] [70] [33] [71] [40], for tree=fill=child-five-color, text width=20em] ] [ Distillation Attack, for tree=fill=child-three-color [ [72] [73] [32] [74] [75], for tree=fill=child-five-color, text width=20em] ] [ Tampering Attack, for tree=fill=child-three-color [ [35] [34] [76] [38] [77] [78] [39], for tree=fill=child-five-color, text width=20em] ] ] [ Inference Stage III-A2, for tree=fill=child-two-color, [ Prompt Attack, for tree=fill=child-three-color [ [29] [27] [79] [37] [80] [36], for tree=fill=child-five-color, text width=20em] ] [ Adversarial Attack, for tree=fill=child-three-color [ [81] [82] [83] [22] [70] [84] [85] [31] [86] [43], for tree=fill=child-five-color, text width=25em] ] [ Jailbreak Chain, for tree=fill=child-three-color [ [87] [25] [26] [30] [28] [88] [89] [90] [41] [91], for tree=fill=child-five-color, text width=25em] ] ] ] [ Hierarchy Based on IoA §$\S$ III-A, for tree=fill=child-one-color, [ Prompt Level III-A3, for tree=fill=child-two-color, [ Prompt Disguise, for tree=fill=child-three-color, [Completion, for tree=fill=child-four-color, text width=15em, [ [89], for tree=fill=child-five-color, text width=20em ] ] [Replace, for tree=fill=child-four-color, text width=15em [ [85] [27] [90], for tree=fill=child-five-color, text width=20em ] ] [Low-resource Language, for tree=fill=child-four-color, text width=15em [ [90], for tree=fill=child-five-color, text width=20em ] ] [Multi-strategy Fusion, for tree=fill=child-four-color, text width=15em [ [31] [33] [90] [88] [28] [43] [71] [75], for tree=fill=child-five-color, text width=20em ] ] ] [ Prompt Rewrite, for tree=fill=child-three-color [Multi-query Optimization, for tree=fill=child-four-color, text width=15em [ [25] [26] [41] [43] [22] [85] [32] [33] [30] [90] [79] [29] [89] [91] [80] [84], for tree=fill=child-five-color, text width=32em ] ] ] ] [ Inference Level III-A4, for tree=fill=child-two-color, [ Scene Nesting, for tree=fill=child-three-color [Story, for tree=fill=child-four-color, text width=15em [ [26] [36] [79] [82] [37] [40] [86] [74] [91], for tree=fill=child-five-color, text width=20em ] ] [Code, for tree=fill=child-four-color, text width=15em [ [26] [37], for tree=fill=child-five-color, text width=20em ] ] [Table, for tree=fill=child-four-color, text width=15em [ [26], for tree=fill=child-five-color, text width=20em ] ] ] [ RAG, for tree=fill=child-three-color [ [77] [78] [39] , for tree=fill=child-five-color, text width=20em ] ] ] [ Model Level III-A5, for tree=fill=child-two-color, [ Gradient-based, for tree=fill=child-three-color [Gradient-based ASG, for tree=fill=child-four-color, text width=15em [ [41] [81] [32] [29] [82] [69] [70] [68] [43], for tree=fill=child-five-color, text width=20em ] ] [Data Poisoning, for tree=fill=child-four-color, text width=15em [ [35] [38] [73] [34] [76], for tree=fill=child-five-color, text width=20em ] ] ] [ Fine-tuning Attack, for tree=fill=child-three-color [ [87] [72] [83], for tree=fill=child-five-color, text width=20em ] ] ] ] ]
Figure 5: The classification of existing universal jailbreak methods. We categorize the jailbreak methods into ”Stages Based on the Impact of Attacks (Stages Based on IoA)” and ”Hierarchy Based on the Impact of Attacks (Hierarchy Based on IoA)”.
@@
As illustrated in Figure 5, impact of attack can be divided into two categories: Stages Based on the Impact of Attacks and Hierarchy Based on the Impact of Attacks. The former primarily targets the inference stage and the training stage, while the latter pertains to different hierarchical levels, including the prompt level, inference level, and model level.
• Stages Based on the Impact of Attacks

### v3

@@ Abstract
Large language models (LLMs) are rapidly evolving from single-modal systems to multimodal LLMs and intelligent agents, significantly expanding their capabilities while introducing increasingly severe security risks. This paper presents a systematic survey of the growing complexity of jailbreak attacks and corresponding defense mechanisms within the expanding LLM ecosystem. We first trace the developmental trajectory from LLMs to MLLMs and Agents, highlighting the core security challenges emerging at each stage. Next, we categorize mainstream jailbreak techniques from both the attack impact and attacker permissions perspectives, and provide a comprehensive analysis of representative attack methods, related datasets, and evaluation metrics. On the defense side, we organize existing strategies based on response timing and technical approach, offering a structured understanding of their applicability and implementation. Furthermore, we identify key limitations in existing surveys, such as insufficient attention to agent-specific security issues, the absence of a clear taxonomy for hybrid jailbreak methods, a lack of detailed analysis of experimental setups, and outdated coverage of recent advancements. To address these limitations, we provide an updated synthesis of recent work and introduce future research directions in areas such as dataset construction, evaluation framework optimization, and strategy generalization. Our study seeks to enhance the understanding of jailbreak mechanisms and facilitate the advancement of more resilient and adaptive defense strategies in the context of ever more capable LLMs.
@@
LLMs, MLLMs, agents, jailbreak attack, defense strategy.
@@ I-A Development of LLMs
The evolution of neural network architectures undergoes multiple paradigm shifts. Early sequence modeling primarily relies on Recurrent Neural Networks (RNNs [1]), whose performance is limited by the vanishing gradient problem when modeling long-term dependencies. Long Short-Term Memory networks (LSTMs [2]) alleviate this issue to some extent by introducing various gating mechanisms. However, due to their sequential nature, LSTMs remain computationally inefficient when processing large-scale data [3]. The emergence of the Transformer [4] architecture in 2017 fundamentally changes this landscape. Its self-attention mechanism enables global context modeling and supports parallel computation. Furthermore, the use of residual connections [5] and layer normalization [6] makes it feasible to train ultra-deep networks, thereby laying the groundwork for the development of large-scale models.
The rapid growth in model size follows the scaling laws proposed by OpenAI in 2020, specifically the performance–compute power law [7], which shows that model performance improves with increasing model parameters, data volume, and compute resources. The parameter count increases dramatically in just a few years: from BERT’s [8] 340 million parameters in 2018, to GPT-3’s [9] 175 billion in 2020, and to PaLM’s [10] 540 billion in 2022. The extensions of more modalities also become a key focus. CLIP [11] achieves semantic alignment between images and texts through contrastive learning, while ViT [12] validates the Transformer’s feasibility for vision tasks.
In recent years, large language models (LLMs) enter a new phase marked by emergent intelligence. Once model parameters exceed a certain threshold, capabilities such as chain-of-thought reasoning and in-context learning emerge. Technological pathways also diversify: InstructGPT [13] enhances alignment with human intent through instruction tuning; reinforcement learning from human feedback (RLHF [13]) becomes central to value alignment; and parameter-efficient fine-tuning methods such as LoRA [14] reduce adaptation costs. Multimodal integration is also accelerating, with GPT-4V [15] supporting both visual and textual understanding and generation. Alongside these trends, architectural innovations continue to reshape LLM development. DeepSeek-R1 [16], for instance, builds upon the Mixture of Experts (MoE) paradigm [17, 18], incorporating dynamic expert routing and a shared attention backbone. This design enables efficient expert utilization while maintaining strong generalization across tasks.
@@ I-B Advances in LLM Jailbreaking and Defense
With the rapid development of LLMs, their application scenarios expand from pure text processing to multimodal understanding and autonomous agents, reflecting an evolutionary trend from LLMs to Multimodal LLMs (MLLMs) and further to intelligent agents [19, 20]. This evolution greatly expands the capabilities of such models, enabling them to handle more complex tasks, but it also introduces more severe security challenges [21, 22]. Among these, jailbreak attacks aim at bypassing safety mechanisms and inducing models to produce inappropriate or restricted content, and they become increasingly complex and diverse [23, 24].
@@
Nevertheless, current jailbreak and defense surveys in LLM ecosystems face several limitations: (1) Insufficient focus on agents: Despite the rapid advancement of agent technology, current jailbreak studies predominantly target traditional LLMs. There is a lack of systematic analysis on jailbreak attacks, adversarial strategies, and defenses specific to agents, which hinders a comprehensive understanding of agent security and limits the optimization of defense mechanisms. (2) Inadequate taxonomy of hybrid jailbreak methods: Modern jailbreak techniques evolve into hybrid forms that combine multiple strategies. These often share common core modules, requiring researchers to carefully classify and analyze them to reveal their structural components and execution logic. (3) Lack of detailed analysis of experimental settings: While various datasets and evaluation metrics are used in jailbreak research, existing surveys rarely provide comprehensive mapping between methods and their evaluation frameworks, impeding the comparability of different approaches and obscuring their respective strengths and weaknesses. (4) Difficulty in covering the latest advancements: Due to the rapid progress of jailbreak and defense research, existing reviews may not incorporate the most recent techniques, resulting in outdated insights into current trends.
To address these limitations and enhance the understanding of jailbreak and defense across the LLM ecosystem, a more comprehensive and up-to-date survey is essential. In this study, we systematically review the evolution from LLMs to MLLMs and then to agents, analyzing core technologies, key features, and security challenges at each stage. We also categorize mainstream jailbreak methods from the perspectives of attack impact and attacker permissions, along with a detailed summary of associated datasets and evaluation metrics. Moreover, we classify existing defense strategies by their response timing and technical approaches. Finally, we introduce open research problems, including dataset construction, diversified attack strategies, and evaluation framework optimization, aiming to provide valuable guidance for future research. Figure 1 illustrates the overall framework of this work, which facilitates a clearer understanding of current research and inform future developments in this field.
Figure 1: The overall architectural diagram of this paper.
@@
(1) We provide a comprehensive review of the evolution from LLMs to MLLMs and then to agents, highlighting the task definitions and key features of jailbreak attacks at each stage (Chapter II). By analyzing capability improvements, expanded applications, and emerging security challenges, we offer a systematic background for understanding jailbreak and defense mechanisms.
(2) We categorize jailbreak techniques from two perspectives: the impact of attack and the permissions of attacker (Chapter III). From the perspective of attack impact, we classify existing methods based on the impact stage and the intervention level (Section III-A). From the perspective of attacker permissions, we divide attacks into black-box and white-box categories, and further organize them according to their targets (Section III-B).
@@
(4) We classify existing defense strategies according to response timing and technical approach (Chapter VI). The response timing includes input-level, output-level, and joint defenses (Section VI-A), while technical approaches are grouped into rule/heuristic-based, Machine Learning(ML)/Deep Learning(DL)-based, adversarial detection, and hybrid strategies (Section VI-B).
@@
Figure 2: Attack workflow of traditional LLMs under the text modality. Attacker strategically design adversarial inputs in the textual modality to elicit harmful responses from the model.
Figure 3: Attack workflow of MLLMs. Attacker crafts adversarial inputs by exploiting the vulnerabilities across different modalities in combination, aiming to manipulate the model into generating harmful outputs.
@@ II-A From LLMs to MLLMs to Agents
LLMs are primarily trained on massive text corpora and focus on capabilities in text generation and comprehension. Through self-attention mechanisms, they capture linguistic patterns and can perform tasks such as question answering, writing, and translation [55]. Despite their impressive performance in text-based interactions, LLMs may still produce erroneous or inappropriate content due to biases in training data or prompt manipulation. Therefore, techniques like safety filtering and RLHF [13] are necessary to ensure security. However, LLMs are limited to pure text interaction and lack the ability to process multimodal information such as images or audio.
MLLMs overcome the limitations of single-modality text input by integrating visual, auditory, and other types of data, enabling cross-modal reasoning and generation (e.g., storytelling based on images, summarizing video content) [11]. Their core technologies include cross-modal alignment and joint representation learning, which equip them with richer perceptual and expressive capabilities [56]. However, they still face challenges such as noise in multimodal data alignment, implicit semantic conflicts, and increased computational demands. In addition, MLLMs still rely on human instructions for decision-making and lack autonomous action capabilities.
Compare to LLMs and MLLMs, agents are intelligent entities capable of perceiving their environment and autonomously taking actions to achieve specific goals. They are typically composed of four key components: core, planning, tools, and memory, with each part working in synergy to support the agent’s autonomous decision-making ability. The core of the agent is the LLM, which is responsible for generating outputs and guiding the decision-making process. The tools section includes external applications and software interfaces that the LLM can invoke, such as internet search, database retrieval, and external system control, enabling the agent to be more flexible and efficient when performing tasks. The planning component aims to reduce hallucinations and inaccuracies in the LLM’s outputs by using structured prompts and introducing additional logical frameworks, ensuring that the core model can make more precise decisions when faced with complex tasks. The memory component is another crucial part of the LLM agent, designed to overcome the current context length limitations of LLMs. By effectively managing and storing historical information, it allows the agent to update and retrieve relevant data during ongoing interactions.
@@ II-B Task Definition
As illustrated in Figure 2, jailbreak attacks on LLMs represent a security threat targeting advanced natural language processing systems. Attackers craft prompts or input sequences to bypass the model’s safety mechanisms, thereby inducing the model to generate content that violates ethical guidelines or contains harmful information [57]. Such attacks can be formalized as an optimization problem: by setting a target output $y_{attack}$, the attacker seeks the optimal input $x^{*}$ such that the model’s output is maximized to $y_{attack}$ [58]:
@@
where $M$ denotes the target LLM and $P$ represents the model’s probability distribution function.
@@
${x_{multi}^{*}}=\underset{{x_{multi}}}{\text{argmax}}\,P(M({x_{multi}})=y_{attack}),$ (2)
where ${x_{multi}}$ = $\{x_{text},x_{image},x_{video}\}$ represents the multimodal input, and $M$ is the target MLLM. This type of attack not only increases stealth but also raises the complexity of defense mechanisms [61].
As depicted in Figure 4, jailbreak attacks on agents exhibit distinct characteristics. Their core objective is to alter the agent’s decision-making behavior, causing it to deviate from its predefined objective task [62, 63]. From the perspective of reinforcement learning, such attacks can be executed by manipulating the reward function $R(s)$. Specifically, the attacker seeks an optimal policy $\pi^{*}$ such that:
@@
Although the aforementioned three types of jailbreak attacks differ in terms of targets, implementation methods and attack intentions, they all follow a common pattern of achieving adversarial goals through specific inputs or environmental perturbations. From the perspective of technical complexity, LLM jailbreak attacks primarily focus on the generation of adversarial examples at the textual level [64], MLLM jailbreaks involve joint optimization of multimodal input [65], while attacks on agents require a deep understanding of the agent’s decision-making mechanisms, task planning, and execution frameworks [66]. The evolution of these attack methods not only reflects the increasing complexity of security threats to AI systems but also underscores the importance of systematically organizing and categorizing these attack strategies.
@@ III Jailbreak Methods
In this section, we focus on various jailbreak attack methods from two distinct perspectives: the attack impact perspective and the attacker permissions perspective. We begin with a conceptual overview from the perspective of attack impact, introducing generalizable modules that are applicable across various jailbreak methods. Then, we conduct an in-depth analysis of specific methods proposed by different researchers based on the attacker’s access permissions.
@@ III-A Impact of Attack
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=root-line-color, rounded corners, align=center, text centered, minimum width=5em, edge path= [draw=edge-color, line width=1pt] (!parent.east) – +(5pt,0) —- (.child anchor); , s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1pt, par/.style=rotate=0, child anchor=north, parent anchor=south, anchor=center, minimum height=3em, , where level=1 draw=child-one-line-color, text width=18em, minimum height=3em, font=, , where level=2 draw=child-two-line-color, text width=18em, minimum height=3em, font=, , where level=3 draw=child-three-line-color, minimum width=18em, minimum height=3em, font=, , where level=4 draw=child-four-line-color, minimum height=3em, font=, , where level=5 draw=child-five-line-color, minimum height=3em, font=, , [ Impact of Attack, par, for tree=fill=root-color, [ Impact Stage of Attack $\S$ III-A, for tree=fill=child-one-color, [ Training Stage III-A1, for tree=fill=child-two-color, [ Backdoor Attack, for tree=fill=child-three-color [ [67] [68] [69] [33] [70] [40], for tree=fill=child-five-color, text width=20em] ] [ Distillation Attack, for tree=fill=child-three-color [ [71] [72] [32] [73] [74], for tree=fill=child-five-color, text width=20em] ] [ Tampering Attack, for tree=fill=child-three-color [ [35] [34] [75] [38] [76] [77] [39], for tree=fill=child-five-color, text width=20em] ] ] [ Inference Stage III-A2, for tree=fill=child-two-color, [ Prompt Attack, for tree=fill=child-three-color [ [29] [27] [78] [37] [79] [36], for tree=fill=child-five-color, text width=20em] ] [ Adversarial Attack, for tree=fill=child-three-color [ [80] [81] [82] [22] [69] [83] [84] [31] [85] [43], for tree=fill=child-five-color, text width=25em] ] [ Jailbreak Chain, for tree=fill=child-three-color [ [86] [25] [26] [30] [28] [87] [88] [89] [41] [90], for tree=fill=child-five-color, text width=25em] ] ] ] [ Intervention Level of Attack $\S$ III-A, for tree=fill=child-one-color, [ Prompt Level III-A3, for tree=fill=child-two-color, [ Prompt Disguise, for tree=fill=child-three-color, [Completion, for tree=fill=child-four-color, text width=15em, [ [88], for tree=fill=child-five-color, text width=20em ] ] [Replace, for tree=fill=child-four-color, text width=15em [ [84] [27] [89], for tree=fill=child-five-color, text width=20em ] ] [Low-resource Language, for tree=fill=child-four-color, text width=15em [ [89], for tree=fill=child-five-color, text width=20em ] ] [Multi-strategy Fusion, for tree=fill=child-four-color, text width=15em [ [31] [33] [89] [87] [28] [43] [70] [74], for tree=fill=child-five-color, text width=20em ] ] ] [ Prompt Rewrite, for tree=fill=child-three-color [Multi-query Optimization, for tree=fill=child-four-color, text width=15em [ [25] [26] [41] [43] [22] [84] [32] [33] [30] [89] [78] [29] [88] [90] [79] [83], for tree=fill=child-five-color, text width=32em ] ] ] ] [ Inference Level III-A4, for tree=fill=child-two-color, [ Scene Nesting, for tree=fill=child-three-color [Story, for tree=fill=child-four-color, text width=15em [ [26] [36] [78] [81] [37] [40] [85] [73] [90], for tree=fill=child-five-color, text width=20em ] ] [Code, for tree=fill=child-four-color, text width=15em [ [26] [37], for tree=fill=child-five-color, text width=20em ] ] [Table, for tree=fill=child-four-color, text width=15em [ [26], for tree=fill=child-five-color, text width=20em ] ] ] [ RAG, for tree=fill=child-three-color [ [76] [77] [39] , for tree=fill=child-five-color, text width=20em ] ] ] [ Model Level III-A5, for tree=fill=child-two-color, [ Gradient-based, for tree=fill=child-three-color [Gradient-based ASG, for tree=fill=child-four-color, text width=15em [ [41] [80] [32] [29] [81] [68] [69] [67] [43], for tree=fill=child-five-color, text width=20em ] ] [Data Poisoning, for tree=fill=child-four-color, text width=15em [ [35] [38] [72] [34] [75], for tree=fill=child-five-color, text width=20em ] ] ] [ Fine-tuning Attack, for tree=fill=child-three-color [ [86] [71] [82], for tree=fill=child-five-color, text width=20em ] ] ] ] ]
Figure 5: The categorization of existing jailbreak methods from the perspective of attack impact. We categorize the jailbreak methods into ”Impact Stage of Attack” and ”Intervention Level of Attack”.
@@
As illustrated in Figure 5, the impact of jailbreak attacks can be categorized from the following two perspectives: Impact Stage of Attack and Intervention Level of Attack. The former primarily targets the inference stage and the training stage, while the latter pertains to different hierarchical levels, including the prompt level, inference level, and model level.
• Impact Stage of Attack